OpenStudy (sophiesky):

Solve the equation. I need help understanding how to do this for my precalc review. Had this question up earlier but I put the wrong equation and think it was confusing people. Posting the equation below on comments. Thank you!

OpenStudy (sophiesky):

\[\frac{ -3x }{ x^2-4x-32 }-\frac{ 2 }{ x-8 }=\frac{ 3 }{ x+4 }\]

OpenStudy (sophiesky):

@pooja195

OpenStudy (sunnnystrong):

Step one: Find the LCD & Multiply throughout

OpenStudy (sunnnystrong):

@sophiesky .. How would you factor \[x^2-4x-32\]

OpenStudy (sophiesky):

Multiply both sides by that?

OpenStudy (sunnnystrong):

\[\frac{ -3x }{ (x-8)(x+4) }-\frac{ 2 }{ (x-8) }=\frac{ 3 }{ (x+4) }\] ** After you factor... the lcd is (x-8)(x+4)...

OpenStudy (sophiesky):

Wait so how did you do that?

OpenStudy (sunnnystrong):

\[x^2-4x-32\] * Factor: What factors of 32 when multiplied together = (-32) & when added together = (-4) -8 & 4 so: (x-8)(x+4)

OpenStudy (sophiesky):

Ohhhhh okay I'm following you know. :)

OpenStudy (sunnnystrong):

Yep! So one you do that... You will multiply throughout.. Like this: \[\frac{ -3x }{ (x-8)(x+4) }-\frac{ 2(x+4) }{ (x-8)(x+4) }=\frac{ 3(x-8) }{ (x+4)(x-8) }\] To get LCD ---> \[-3x-2(x+4)=3(x-8)\] You are left with this ^

OpenStudy (sophiesky):

Alright, and I continue to work it out right? Like that isn't the finished product?

OpenStudy (sophiesky):

Okay x is 2!

OpenStudy (sophiesky):

Mind helping me with another @sunnnystrong ? :)